# Pretraining Decision Transformers with Reward Prediction for In-Context Multi-task Structured Bandit Learning

### v2

@@ A.14 Empirical Study: Offline Performance
We first discuss how PreDeToR (-$\tau$) is modified for the offline setting. In the offline setting, the PreDeToR first samples a task $m\sim\mathcal{T}_{\text{test}}$, then the test dataset $\mathcal{H}_{m}\sim\mathcal{D}_{\text{test}}(\cdot|m)$. Then PreDeToR and PreDeToR-$\tau$ act similarly to the online setting, but based on the entire offline dataset $\mathcal{H}_{m}$. The full pseudocode of PreDeToR is in Algorithm 2.
@@
Recall that $\mathcal{D}_{\text{test}}$ denote a distribution over all possible interactions that can be generated by $\pi^{w}$ during test time. For offline testing, first, a test task $m\sim\mathcal{T}_{\text{test}}$, and then an in-context test dataset $\mathcal{H}_{m}$ is collected such that $\mathcal{H}_{m}\sim\mathcal{D}_{\text{test}}(\cdot|m)$. Observe from Algorithm 2 that in the offline setting, PreDeToR first samples a task $m\sim\mathcal{T}_{\text{test}}$, and then a test dataset $\mathcal{H}_{m}\sim\mathcal{D}_{\text{test}}(\cdot|m)$ and acts greedily. Crucially in the offline setting the PreDeToR does not add the observed reward $r_{t}$ at round $t$ to the dataset. Through this experiment, we want to evaluate the performance of PreDeToR to learn the underlying latent structure and reward correlation when the horizon is small. Finally, recall that when the horizon is small the weak demonstrator $\pi^{w}$ does not have sufficient samples for each action. This leads to a poor approximation of the greedy action.
Baselines: We again implement the same baselines discussed in Section 4. The baselines are PreDeToR, PreDeToR-$\tau$, DPT-greedy, AD, Thomp, and LinUCB. During test time evaluation for offline setting the DPT selects $I_{t}=\widehat{a}_{m,t,*}$ where $\widehat{a}_{m,t,*}=\mathop{\rm arg\,max}_{a}\mathrm{TF}_{\mathbf{\Theta}}(a|\mathcal{H}^{t}_{m})$ is the predicted optimal action.
@@
Experimental Result: We observe these outcomes in Figure 21. In Figure 21 we show the linear bandit setting for horizon $n=20$, $M_{\text{pre}}=200000$, $M_{\text{test}}=5000$, $A=20$, and $d=5$ for the low data regime. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has comparable cumulative regret to DPT-greedy. Note that for any task $m$ for the horizon $n=20$ the Thomp will be able to sample all the actions at most once. In the non-linear setting of Figure 21(b) the $n=40$, $M_{\text{pre}}=100000$, $A=6$, $d=2$. Observe that in all of these results, the performance of PreDeToR (-$\tau$) is comparable with respect to cumulative regret against DPT-greedy.
@@ B.1 Proof of 1
The B.1 ensures that the matrix $\left(\sigma^{2}_{\bm{\theta}}\mathbf{H}_{n}(\mathbf{S}_{A}+\mathbf{D}_{A})\mathbf{H}_{n}^{\top}\right)^{-1}$ is invertible. Under B.1, plugging the estimate $\mathbf{S}_{A}$ back in (8) shows that the average posterior mean over all the tasks is
@@ C.1 Generalization Proof
(PreDeToR risk) Suppose error stability 8.1 holds and assume loss function $\ell(\cdot,\cdot)$ is $C$-Lipschitz for all $r_{t}\in[0,B]$ and horizon $n\geq 1$. Let $\widehat{\mathrm{TF}}$ be the empirical solution of (ERM) and $\mathcal{N}(\mathcal{A},\rho,\epsilon)$ be the covering number of the algorithm space $\mathrm{Alg}$ following Definition C.2 and C.3. Then with probability at least $1-2\delta$, the excess Multi-task learning (MTL) risk of PreDeToR-$\tau$ is bounded by
@@
^$\widehat{\mathcal{L}}_{\mathcal{H}_{\mathrm{all}}}(\mathrm{TF})=\frac{1}{nM}\sum_{m=1}^{M}\sum_{t=1}^{n}\ell\left(r_{mt}(a_{mt}),\mathrm{TF}\left(\widehat{r}_{mt}(a_{mt})|\mathcal{H}_{m}^{t-1},a_{mt}\right)\right)$
@@
In the following discussion, we drop the subscripts MTL and $\mathcal{H}_{\text{all. }}$ The excess MTL risk is decomposed as follows:
$\mathcal{R}_{\mathrm{MTL}}(\widehat{\mathrm{TF}})$ $=\mathcal{L}(\widehat{\mathrm{TF}})-\mathcal{L}\left(\mathrm{TF}^{*}\right)$
@@
^$\widehat{\mathcal{L}}_{m}(\mathrm{TF}):=\frac{1}{n}\sum_{t=1}^{n}\ell\left(r_{mt}(a_{mt}),\mathrm{TF}\left(\widehat{r}_{mt}(a_{mt})|\mathcal{H}_{m}^{t-1},a_{mt}\right)\right),\quad\text{ and }$
@@
Define the random variables $X_{m,t}=\mathbb{E}\left[\widehat{\mathcal{L}}_{t}(\mathrm{TF})\mid\mathcal{H}_{m}^{t}\right]$ for $t\in[n]$ and $m\in[M]$, that is, $X_{m,t}$ is the expectation over ^$\widehat{\mathcal{L}}_{t}(\mathrm{TF})$ given training sequence $\mathcal{H}_{m}^{t}=\left\{\left(a_{mt^{\prime}},r_{mt^{\prime}}\right)\right\}_{t^{\prime}=1}^{t}$ (which are the filtrations). With this, we have that $X_{m,n}=\mathbb{E}\left[\widehat{\mathcal{L}}_{m}(\mathrm{TF})\mid\mathcal{H}_{m}^{n}\right]=\widehat{\mathcal{L}}_{m}(\mathrm{TF})$ and $X_{m,0}=\mathbb{E}\left[\widehat{\mathcal{L}}_{m}(\mathrm{TF})\right]=\mathcal{L}_{m}(\mathrm{TF})$. More generally, $\left(X_{m,0},X_{m,1},\ldots,X_{m,n}\right)$ is a martingale sequence since, for every $m\in[M]$, we have that $\mathbb{E}\left[X_{m,i}\mid\mathcal{H}_{m}^{t-1}\right]=X_{m,t-1}$. For notational simplicity, in the following discussion, we omit the subscript $m$ from $a,r$ and $\mathcal{H}$ as they will be clear from the left-hand-side variable $X_{m,t}$. We have that
@@
|^$|\widehat{\mathcal{L}}(\mathrm{TF})-\mathcal{L}(\mathrm{TF})|\leq(B+K\log n)\sqrt{\frac{\log(1/\delta)}{cnM}}$ (11)
@@
and we aim to bound $\sup_{\mathrm{TF}\in\mathrm{Alg}}|h(\mathrm{TF})|$. Following Definition C.3, for $\varepsilon>0$, let $\mathrm{Alg}_{\varepsilon}$ be a minimal $\varepsilon$-cover of $\mathrm{Alg}$ in terms of distance metric $\rho$. Therefore, $\mathrm{Alg}_{\varepsilon}$ is a discrete set with cardinality $\left|\mathrm{Alg}_{\varepsilon}\right|:=\mathcal{N}(\mathrm{Alg},\rho,\varepsilon)$. Then, we have
@@
We will first bound the quantity $\sup_{\mathrm{TF}\in\mathrm{Alg}^{\prime}}\min_{\mathrm{TF}\in\mathrm{Alg}_{\varepsilon}}\left|h(\mathrm{TF})-h\left(\mathrm{TF}{}^{\prime}\right)\right|$. We will utilize that loss function $\ell(\cdot,\cdot)$ is $C$-Lipschitz. For any $\mathrm{TF}\in\mathrm{Alg}$, let $\mathrm{TF}\in\mathrm{Alg}_{\varepsilon}$ be its neighbor following Definition C.3. Then we can show that
|^(TF)^$\left|\widehat{\mathcal{L}}(\mathrm{TF})-\widehat{\mathcal{L}}\left(\mathrm{TF}^{\prime}\right)\right|$
@@
(Stability Factor) The work of Li et al. (2023) also characterizes the stability factor $K$ in 8.1 with respect to the transformer architecture. Assuming loss $\ell(\cdot,\cdot)$ is C-Lipschitz, the algorithm induced by $\mathrm{TF}(\cdot)$ obeys the stability assumption with $K=2C\left((1+\Gamma)e^{\Gamma}\right)^{L}$, where the norm of the transformer weights are upper bounded by $O(\Gamma)$ and there are $L$-layers of the transformer.
@@ C.2 Generalization Error to New Task
[…]
@@
Let the target task $g$ be sampled from $\mathcal{T}$, and the test set $\mathcal{H}_{g}=\{a_{t},r_{t}\}_{t=1}^{n}$. Define empirical and population risks on $g$ as ^$\widehat{\mathcal{L}}_{g}(\mathrm{TF})=\frac{1}{n}\sum_{t=1}^{n}\ell\left(r_{t}(a_{mt}),\mathrm{TF}\left(\widehat{r}_{t}(a_{mt})|\mathcal{H}_{g}^{t-1},a_{t}\right)\right)$ and $\mathcal{L}_{g}(\mathrm{TF})=\mathbb{E}_{\mathcal{H}_{g}}\left[\widehat{\mathcal{L}}_{g}(\mathrm{TF})\right]$. Again we drop $\mathbf{\Theta}$ from the transformer notation. Then the expected excess transfer risk following (ERM) is defined as
@@
^$\widehat{\mathcal{L}}_{m}(\mathrm{TF}):=\frac{1}{n}\sum_{t=1}^{n}\ell\left(r_{mt}(a_{mt}),\mathrm{TF}\left(\widehat{r}_{mt}(a_{mt})|\mathcal{D}_{m}^{t-1},a_{mt}\right)\right),\quad\text{ and }$
@@
$\mathbb{E}_{g}\left[\mathcal{R}_{g}(\widehat{\mathrm{TF}})\right]$ $=\mathbb{E}_{g}\left[\mathcal{L}_{g}(\widehat{\mathrm{TF}})\right]-\mathbb{E}_{g}\left[\mathcal{L}_{g}\left(\mathrm{TF}^{*}\right)\right]$
@@
Next, let $\mathrm{Alg}_{\varepsilon}$ be the minimal $\varepsilon$-cover of $\mathrm{Alg}$ following Definition C.2, which implies that for any task $g\sim\mathcal{T}$, and any $\mathrm{TF}\in\mathrm{Alg}$, there exists $\mathrm{TF}^{\prime}\in\mathrm{Alg}_{\varepsilon}$

### v3

@@ A.14 Empirical Study: Offline Performance
We first discuss how PreDeToR (-$\tau$) is modified for the offline setting. In the offline setting, the PreDeToR first samples a task $m\sim\mathcal{T}_{\text{test}}$, then the test dataset $\mathcal{H}_{m}\sim\mathcal{D}_{\text{test}}(\cdot|m)$. Then PreDeToR and PreDeToR-$\tau$ act similarly to the online setting, but based on the entire offline dataset $\mathcal{H}_{m}$. The full pseudocode of PreDeToR is in Algorithm˜2.
@@
Recall that $\mathcal{D}_{\text{test}}$ denote a distribution over all possible interactions that can be generated by $\pi^{w}$ during test time. For offline testing, first, a test task $m\sim\mathcal{T}_{\text{test}}$, and then an in-context test dataset $\mathcal{H}_{m}$ is collected such that $\mathcal{H}_{m}\sim\mathcal{D}_{\text{test}}(\cdot|m)$. Observe from Algorithm˜2 that in the offline setting, PreDeToR first samples a task $m\sim\mathcal{T}_{\text{test}}$, and then a test dataset $\mathcal{H}_{m}\sim\mathcal{D}_{\text{test}}(\cdot|m)$ and acts greedily. Crucially in the offline setting the PreDeToR does not add the observed reward $r_{t}$ at round $t$ to the dataset. Through this experiment, we want to evaluate the performance of PreDeToR to learn the underlying latent structure and reward correlation when the horizon is small. Finally, recall that when the horizon is small the weak demonstrator $\pi^{w}$ does not have sufficient samples for each action. This leads to a poor approximation of the greedy action.
Baselines: We again implement the same baselines discussed in Section˜4. The baselines are PreDeToR, PreDeToR-$\tau$, DPT-greedy, AD, Thomp, and LinUCB. During test time evaluation for offline setting the DPT selects $I_{t}=\widehat{a}_{m,t,*}$ where $\widehat{a}_{m,t,*}=\mathop{\rm arg\,max}_{a}\mathrm{TF}_{\mathbf{\Theta}}(a|\mathcal{H}^{t}_{m})$ is the predicted optimal action.
@@
Experimental Result: We observe these outcomes in Figure˜21. In Figure˜21 we show the linear bandit setting for horizon $n=20$, $M_{\text{pre}}=200000$, $M_{\text{test}}=5000$, $A=20$, and $d=5$ for the low data regime. Again, the demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has comparable cumulative regret to DPT-greedy. Note that for any task $m$ for the horizon $n=20$ the Thomp will be able to sample all the actions at most once. In the non-linear setting of Figure˜21(b) the $n=40$, $M_{\text{pre}}=100000$, $A=6$, $d=2$. Observe that in all of these results, the performance of PreDeToR (-$\tau$) is comparable with respect to cumulative regret against DPT-greedy.
@@ B.1 Proof of ˜1
The ˜B.1 ensures that the matrix $\left(\sigma^{2}_{\bm{\theta}}\mathbf{H}_{n}(\mathbf{S}_{A}+\mathbf{D}_{A})\mathbf{H}_{n}^{\top}\right)^{-1}$ is invertible. Under ˜B.1, plugging the estimate $\mathbf{S}_{A}$ back in (8) shows that the average posterior mean over all the tasks is
@@ C.1 Generalization Proof
(PreDeToR risk) Suppose error stability ˜8.1 holds and assume loss function $\ell(\cdot,\cdot)$ is $C$-Lipschitz for all $r_{t}\in[0,B]$ and horizon $n\geq 1$. Let $\widehat{\mathrm{TF}}$ be the empirical solution of (ERM) and $\mathcal{N}(\mathcal{A},\rho,\epsilon)$ be the covering number of the algorithm space $\mathrm{Alg}$ following Definition C.2 and C.3. Then with probability at least $1-2\delta$, the excess Multi-task learning (MTL) risk of PreDeToR-$\tau$ is bounded by
@@
$\widehat{\mathcal{L}}_{\mathcal{H}_{\mathrm{all}}}(\mathrm{TF})=\frac{1}{nM}\sum_{m=1}^{M}\sum_{t=1}^{n}\ell\left(r_{mt}(a_{mt}),\mathrm{TF}\left(\widehat{r}_{mt}(a_{mt})|\mathcal{H}_{m}^{t-1},a_{mt}\right)\right)$
@@
[…]
@@
$\widehat{\mathcal{L}}_{m}(\mathrm{TF}):=\frac{1}{n}\sum_{t=1}^{n}\ell\left(r_{mt}(a_{mt}),\mathrm{TF}\left(\widehat{r}_{mt}(a_{mt})|\mathcal{H}_{m}^{t-1},a_{mt}\right)\right),\quad\text{ and }$
@@
Define the random variables $X_{m,t}=\mathbb{E}\left[\widehat{\mathcal{L}}_{t}(\mathrm{TF})\mid\mathcal{H}_{m}^{t}\right]$ for $t\in[n]$ and $m\in[M]$, that is, $X_{m,t}$ is the expectation over $\widehat{\mathcal{L}}_{t}(\mathrm{TF})$ given training sequence $\mathcal{H}_{m}^{t}=\left\{\left(a_{mt^{\prime}},r_{mt^{\prime}}\right)\right\}_{t^{\prime}=1}^{t}$ (which are the filtrations). With this, we have that $X_{m,n}=\mathbb{E}\left[\widehat{\mathcal{L}}_{m}(\mathrm{TF})\mid\mathcal{H}_{m}^{n}\right]=\widehat{\mathcal{L}}_{m}(\mathrm{TF})$ and $X_{m,0}=\mathbb{E}\left[\widehat{\mathcal{L}}_{m}(\mathrm{TF})\right]=\mathcal{L}_{m}(\mathrm{TF})$. More generally, $\left(X_{m,0},X_{m,1},\ldots,X_{m,n}\right)$ is a martingale sequence since, for every $m\in[M]$, we have that $\mathbb{E}\left[X_{m,i}\mid\mathcal{H}_{m}^{t-1}\right]=X_{m,t-1}$. For notational simplicity, in the following discussion, we omit the subscript $m$ from $a,r$ and $\mathcal{H}$ as they will be clear from the left-hand-side variable $X_{m,t}$. We have that
@@
$|\widehat{\mathcal{L}}(\mathrm{TF})-\mathcal{L}(\mathrm{TF})|\leq(B+K\log n)\sqrt{\frac{\log(1/\delta)}{cnM}}$ (11)
@@
and we aim to bound $\sup_{\mathrm{TF}\in\mathrm{Alg}}|h(\mathrm{TF})|$. Following Definition˜C.3, for $\varepsilon>0$, let $\mathrm{Alg}_{\varepsilon}$ be a minimal $\varepsilon$-cover of $\mathrm{Alg}$ in terms of distance metric $\rho$. Therefore, $\mathrm{Alg}_{\varepsilon}$ is a discrete set with cardinality $\left|\mathrm{Alg}_{\varepsilon}\right|:=\mathcal{N}(\mathrm{Alg},\rho,\varepsilon)$. Then, we have
@@
We will first bound the quantity $\sup_{\mathrm{TF}\in\mathrm{Alg}^{\prime}}\min_{\mathrm{TF}\in\mathrm{Alg}_{\varepsilon}}\left|h(\mathrm{TF})-h\left(\mathrm{TF}{}^{\prime}\right)\right|$. We will utilize that loss function $\ell(\cdot,\cdot)$ is $C$-Lipschitz. For any $\mathrm{TF}\in\mathrm{Alg}$, let $\mathrm{TF}\in\mathrm{Alg}_{\varepsilon}$ be its neighbor following Definition˜C.3. Then we can show that
$\left|\widehat{\mathcal{L}}(\mathrm{TF})-\widehat{\mathcal{L}}\left(\mathrm{TF}^{\prime}\right)\right|$
@@
(Stability Factor) The work of Li et al. (2023) also characterizes the stability factor $K$ in ˜8.1 with respect to the transformer architecture. Assuming loss $\ell(\cdot,\cdot)$ is C-Lipschitz, the algorithm induced by $\mathrm{TF}(\cdot)$ obeys the stability assumption with $K=2C\left((1+\Gamma)e^{\Gamma}\right)^{L}$, where the norm of the transformer weights are upper bounded by $O(\Gamma)$ and there are $L$-layers of the transformer.
@@ C.2 Generalization Error to New Task
(Transfer Risk) Consider the setting of Theorem˜8.2 and assume the source tasks are independently drawn from task distribution $\mathcal{T}$. Let  TF $\widehat$$\text${ TF }} be the empirical solution of (ERM) and $g\sim\mathcal{T}$. Then with probability at least $1-2\delta$, the expected excess transfer learning risk is bounded by
@@
Let the target task $g$ be sampled from $\mathcal{T}$, and the test set $\mathcal{H}_{g}=\{a_{t},r_{t}\}_{t=1}^{n}$. Define empirical and population risks on $g$ as $\widehat{\mathcal{L}}_{g}(\mathrm{TF})=\frac{1}{n}\sum_{t=1}^{n}\ell\left(r_{t}(a_{mt}),\mathrm{TF}\left(\widehat{r}_{t}(a_{mt})|\mathcal{H}_{g}^{t-1},a_{t}\right)\right)$ and $\mathcal{L}_{g}(\mathrm{TF})=\mathbb{E}_{\mathcal{H}_{g}}\left[\widehat{\mathcal{L}}_{g}(\mathrm{TF})\right]$. Again we drop $\mathbf{\Theta}$ from the transformer notation. Then the expected excess transfer risk following (ERM) is defined as
@@
$\widehat{\mathcal{L}}_{m}(\mathrm{TF}):=\frac{1}{n}\sum_{t=1}^{n}\ell\left(r_{mt}(a_{mt}),\mathrm{TF}\left(\widehat{r}_{mt}(a_{mt})|\mathcal{D}_{m}^{t-1},a_{mt}\right)\right),\quad\text{ and }$
@@
[…]
@@
Next, let $\mathrm{Alg}_{\varepsilon}$ be the minimal $\varepsilon$-cover of $\mathrm{Alg}$ following Definition˜C.2, which implies that for any task $g\sim\mathcal{T}$, and any $\mathrm{TF}\in\mathrm{Alg}$, there exists $\mathrm{TF}^{\prime}\in\mathrm{Alg}_{\varepsilon}$